3. Ted wants to do a one-year training that costs 20,000. He will lose 100,000 in income if he does it. He is promised a new job at 130,000 per year for five years. After the fifth year, his training will be obsolete. Ted is going to retire anyway. Assume a 5 percent discount rate. Should he do it?
Answer #1

Ans:

computation of net present value

Net Present value = cash flow * Discounting factor

Discounting factor = 1/(1+r)^n

Sum of discounting factor for year 1 to year 5 = 4.3294

Net Present Value = -120,000 * 1/(1+0.05)^0 + 130,000 * sum of discounting factor for year 1 to year 5

= -120,000 * 1 + 130,000 * 4.3294  

= -120,000 + 562,822

= 442822

Since the net present value is positive, he should do the training.